Give as much information as you can about the P-value of a t test in each of the following situations. (Round your answers to three decimal places.)

(a) Upper-tailed test, df = 9, t = 2.0 P-value < 0.005 0.005 < P-value < 0.01 0.01 < P-value < 0.025 0.025 < P-value < 0.05 P-value > 0.05

(b) Upper-tailed test, n = 13, t = 3.2 P-value < 0.005 0.005 < P-value < 0.01 0.01 < P-value < 0.025 0.025 < P-value < 0.05 P-value > 0.05

(c) Lower-tailed test, df = 10, t = −2.4 P-value < 0.005 0.005 < P-value < 0.01 0.01 < P-value < 0.025 0.025 < P-value < 0.05 P-value > 0.05

(d) Lower-tailed test, n = 23, t = −4.2 P-value < 0.005 0.005 < P-value < 0.01 0.01 < P-value < 0.025 0.025 < P-value < 0.05 P-value > 0.05

(e) Two-tailed test, df = 15, t = −1.7 P-value < 0.01 0.01 < P-value < 0.02 0.02 < P-value < 0.05 0.05 < P-value < 0.1 P-value > 0.1

(f) Two-tailed test, n = 16, t = 1.7 P-value < 0.01 0.01 < P-value < 0.02 0.02 < P-value < 0.05 0.05 < P-value < 0.1 P-value > 0.1

(g) Two-tailed test, n = 15, t = 6.3 P-value < 0.01 0.01 < P-value < 0.02 0.02 < P-value < 0.05 0.05 < P-value < 0.1 P-value > 0.1
Answer #1

Solution:-

(a) Upper-tailed test, df = 9, t = 2.0

p-value = 0.0383

0.0383 < 0.05

0.025 < P-value < 0.05

b)  Upper-tailed test, n = 13, t = 3.2

df = n -1 =13-1 = 12

p-value = 0.0038

0.0038 < 0.005

P-value > 0.05 = 0.0038 > 0.05

c) (c) Lower-tailed test, df = 10, t = −2.4

p-value = 0.0187

0.0187 < 0.05

0.01 < P-value < 0.025

d) Lower-tailed test, n = 23, t = −4.2

df = n -1 =23-1 = 22

p-value = 0.0001

0.0001< 0.005

e) Two-tailed test, df = 15, t = −1.7

p-value = 0.1098

0.1098 > 0.05

f)Two-tailed test, n = 16, t = 1.7

df = n-1 =16-1 =15

p-value = 0.1098

0.1098 > 0.05

f) Two-tailed test, n = 15, t = 6.3

df = n-1 =15-1 =14

p-value = 0.0000

0 < 0.01